A risk-free investment of $6000 will return 3.5%. A risky $6000 investment has a 21% chance of defaulting and returning only $3500. How much must the risky investment promise to return?

a. 11.8%                                                 
b. .9.5%
c. 10.5%
d. 15.5%
Answer #1

Option d

6000*(1+3.5%)=(1-21%)*6000*(1+r)+21%*3500
=>r=(6000*(1+3.5%)-21%*3500)/((1-21%)*6000)-1
=>r=15.50633%